OpenStudy (anonymous):

In linear equations, how do i know if i should subtract from the right side or the left side?

OpenStudy (karatechopper):

example..

OpenStudy (anonymous):

I like my coefficient for x to remain positive; therefore, if I have 2x + 10 = 12x I subtract 2x from the left side and if I have 12x + 10 = 2x I subtract 2x from the right side

OpenStudy (anonymous):

3x + 7 = 10x +5 7 = 7x +5 - ^here? ^ or here?

OpenStudy (turingtest):

it really doesn't matter so long as you keep the goal in mind: get x by itself pick one and we can continue from there

OpenStudy (anonymous):

And how much do i subtract? the whole number or so it remains +/-0?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so i got 0=7x-2

OpenStudy (anonymous):

now i need to divide -2 by 7?

OpenStudy (turingtest):

no, we need the 2 on the other side of the equation that's what I mean about getting x by itself

OpenStudy (anonymous):

Ok so i dont do 0=7x-2?

OpenStudy (turingtest):

if we divide by 7 now we get\[0=x-\frac{2}{7}\]which is okay, we can still solve it from here; but it's a good policy to leave fractions for the end if possible. so instead add 2 to both sides to make x a little bit more alone.

OpenStudy (anonymous):

I dont get it, i get like 0.29..

OpenStudy (turingtest):

leave the answer in fraction form

OpenStudy (anonymous):

Oh i see, so i get 2=7x?

OpenStudy (turingtest):

right, now we divide by 7 to isolate x what is x=? (leave as a fraction)

OpenStudy (anonymous):

I dont know.. 2/7?

OpenStudy (turingtest):

you do know! lol shall we do another?

OpenStudy (phi):

If you want more info, watch this and the next 2 videos http://www.khanacademy.org/video/simple-equations?playlist=Algebra

OpenStudy (turingtest):

that's what he's working from phi

OpenStudy (anonymous):

What?

OpenStudy (turingtest):

Another?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Well ill be damned.. it was right!

OpenStudy (turingtest):

got one you want to post? or should I make one up?

OpenStudy (anonymous):

wait, ill try the next one. 6x+3=2x+7

OpenStudy (turingtest):

just go ahead and see how far you get

OpenStudy (anonymous):

So i subtract 2x from both sides, getting 4x+3=7

OpenStudy (turingtest):

good, then?

OpenStudy (anonymous):

Then i subtract 3 from both sides, getting 4x=4

OpenStudy (turingtest):

good, then?

OpenStudy (anonymous):

And then i divide 4 by 4 and get 1 as the final answer :D

OpenStudy (karatechopper):

:D

OpenStudy (karatechopper):

well taught turing:)

OpenStudy (anonymous):

TT, you are an awesome teacher!

OpenStudy (turingtest):

tadah! nice see how you didn't need to ask me which was better, subtract 7 or 3? you knew by looking that 3 was the smart way to go :) practice practice... thanks for the compliment btw

OpenStudy (anonymous):

So where do i go on from here?

OpenStudy (turingtest):

have you done ALL the exercises from this section already?

OpenStudy (anonymous):

Soon. you mean until i can choose another section right?

OpenStudy (turingtest):

Yes. It is imperative that you do every problem. This needs to become second nature for you when you can look at 5x-3=2x+12 and tell me the answer very quickly, with almost no work shown, then you can continue

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

actually 1/3

OpenStudy (turingtest):

nope, better keep working on it. practice practice... don't try to get ahead of yourself, if you need to write it and go slowly, do it. Don't try to shortcut past that.

OpenStudy (anonymous):

alright. but i couldnt figure the last one out, it was 6x+6=0

OpenStudy (anonymous):

I didnt know what to do there. was it 1 or 0? maybe -6?

OpenStudy (turingtest):

just try the problem and show me your steps you don't need to tell me what each step is, btw, I can see by looking

OpenStudy (anonymous):

it was 9x+9=3x+3 6x+9=3 6x+6=0 6x= -6?

OpenStudy (turingtest):

so now what is the last step to get x all alone?

OpenStudy (karatechopper):

tht was correct now show us your next and final step

OpenStudy (anonymous):

6/-6?

OpenStudy (karatechopper):

perfect

OpenStudy (karatechopper):

now simplify that

OpenStudy (anonymous):

-1?

OpenStudy (karatechopper):

correct

OpenStudy (anonymous):

Yes!

OpenStudy (turingtest):

yes, very good what was the last step? dividing by...?

OpenStudy (karatechopper):

that wasn't too bad now was it!

OpenStudy (anonymous):

The number on the right?

OpenStudy (anonymous):

dividing the quantity of x by the number on the right side of =?

OpenStudy (anonymous):

Other*

OpenStudy (turingtest):

no, that is what I was afraid of... though you got the right answer, the -6 should have stayed on top. 6x=-6 what do we divide by to isolate x?

OpenStudy (anonymous):

-6 with 6?

OpenStudy (turingtest):

it's just one answer: divide by 6

OpenStudy (turingtest):

let's try something a little different... solve xy+7=10 for x

OpenStudy (anonymous):

Ok

OpenStudy (turingtest):

Note that above for 6x=-6 we divided both sides by the number next to the x (i.e. 6), that is what I wanted you to see. We do not divide by the number on the right. You will see why in this next problem

OpenStudy (anonymous):

1.5 or 2 for x

OpenStudy (anonymous):

Ok..

OpenStudy (turingtest):

so now xy+7=10 you got a number answer? what happened to y? please show steps.

OpenStudy (karatechopper):

@turing does he know what he is solving for?

OpenStudy (turingtest):

solve for x thanks karate

OpenStudy (karatechopper):

anytime turing:)

OpenStudy (anonymous):

I thought that x times y would be the answer. since 1.5x2 is 3 (which was needed to make 10) i thought that x would be 1.5 or 2.(the other one being y)

OpenStudy (turingtest):

you are over-thinking Inopeki just take each step one-by-one what is the first step?

OpenStudy (anonymous):

Determening what i need to make to complete 10

OpenStudy (turingtest):

no the first step is the same as the others xy+7=10 subtract 7 from both sides

OpenStudy (turingtest):

what do we have after?

OpenStudy (anonymous):

Ah.. xy=3

OpenStudy (turingtest):

good! now again, we want x by itself. it looks like x is being multiplied by y, so how can we get x by itself?

OpenStudy (anonymous):

Removing y from the equation?

OpenStudy (turingtest):

y cannot leave the equation entirely, we can only shift it to one side. I mentioned that x was being multiplied by y. Do you know the inverse operation of multiplication?

OpenStudy (anonymous):

division?

OpenStudy (turingtest):

right! and in this case what do we need to divide by to get x by itself?

OpenStudy (anonymous):

y with 3?

OpenStudy (turingtest):

close, but what do you mean "divide by y with 3"? that makes no sense again it can only be one answer divide by...?

OpenStudy (turingtest):

just to make it visible again: xy=3 what do we divide by on both sides (one answer to this question)

OpenStudy (anonymous):

Ah, divide x and 3 by y?

OpenStudy (turingtest):

YES!!! and what do we have after? write the expression.

OpenStudy (anonymous):

How can i do that when i dont even know what y is?

OpenStudy (turingtest):

Good question. write the answer first and I will tell you what it means after.

OpenStudy (anonymous):

You mean x/y=3/y?

OpenStudy (turingtest):

close, but on the left x should be by itself, that's why we divided by y right? so it should be...

OpenStudy (anonymous):

I GET IT NOW! because the left side was x times y we needed to isolate x so we divided x with y to reverse the effect. now it is x=3/y

OpenStudy (turingtest):

YES!!! very good! but what does it mean...? as you said y can be any number (except zero, but we'll discuss that later) so we have x=3/y that means we have infinite solutions to this problem. To find what they are plug in various numbers for y. So what is x if y=1 y=3 y=10 ???

OpenStudy (anonymous):

1.5?

OpenStudy (turingtest):

no, its different for each one. Remember that I said there were infinite solutions. just plug in the numbers into the equation there is no one answer for x=3/y so plug in a number for y like y=1 what is x?

OpenStudy (anonymous):

3

OpenStudy (turingtest):

yes! what about when y=3?

OpenStudy (anonymous):

1